OpenStudy (anonymous):

Function Notation help! f(x) = 3x + 2 f(3) Is this how you do it: 3(3)+2 9+2 11

Mehek (mehek14):

correct :)

OpenStudy (anonymous):

Thank you! is this also right: g(x) = x^2 -1 g(5) 5^2 - 1 25 - 1 24

jimthompson5910 (jim_thompson5910):

I would keep the equal sign throughout f(x) = 3x + 2 f(3) = 3(3) + 2 f(3) = 9 + 2 f(3) = 11 but other than that, you have the correct steps and final answer.

Mehek (mehek14):

correct agaun

Mehek (mehek14):

again*

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

also, what does it mean when a question says (example): (f)x = 3x + 2 x when (f)x = 2

jimthompson5910 (jim_thompson5910):

the notation should be f(x) not (f)x

jimthompson5910 (jim_thompson5910):

it sounds like they want you to find the value of x when f(x) = 2 so you'll replace all of `f(x)` with 2 and solve for x

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = 3x+2\] \[\Large {\color{red}{f(x)}} = 3x+2\] \[\Large {\color{red}{2}} = 3x+2 \ ... \ \text{Replace } {\color{red}{f(x)}} \text{ with } {\color{red}{2}}\] now solve for x

OpenStudy (anonymous):

sorry i meant to say: x when f(x) = 2